# The Tournament Theorem of Rédei revisited

### v1

@@ 4 Berge’s Stronger Theorem
The above results of Dirac and Berge together imply the following theorem:
@@
(Berge-Dirac Theorem) Let $G$ be a mixed graph with $n\geq 2$ vertices. Let $\mathbb{P}_{0}$ denote the set of permutations of $G$ without elements from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$, let $\mathbb{P}_{i}$ denote the set of permutations of $G$ without elements from $E_{i}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ $(i\in\{1,2\})$, and let $\mathbb{P}_{3}=\mathbb{P}_{1}\cap\mathbb{P}_{2}$. Then:
$|\mathbb{P}_{0}|+|\mathbb{P}_{1}|+|\mathbb{P}_{2}|+|\mathbb{P}_{3}|\mbox{ is even.}$
We do not know any simpler proof of the Berge-Dirac Theorem than the above combination. Assuming the Berge-Dirac Theorem it follows that Dirac’s Corollary 2 and Berge’s Stronger Theorem are equivalent. Also note that in the Berge-Dirac Theorem we may replace any + by a -, for example
$|\mathbb{P}_{0}|-|\mathbb{P}_{1}|-|\mathbb{P}_{2}|+|\mathbb{P}_{3}|$ is even

### v2

@@ 4 Berge’s Stronger Theorem
The above results of Dirac and Berge may be formulated together as follows.
@@
(Berge-Dirac Theorem) Let $G$ be a mixed graph with $n\geq 2$ vertices. Let $\mathbb{P}_{0}$ denote the set of permutations of $G$ without elements from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$, let $\mathbb{P}_{i}$ denote the set of permutations of $G$ without elements from $E_{i}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ $(i\in\{1,2\})$, and let $\mathbb{P}_{3}=\mathbb{P}_{1}\cap\mathbb{P}_{2}$. Then, either $\mathbb{P}_{0}$, $\mathbb{P}_{1}$, $\mathbb{P}_{2}$, $\mathbb{P}_{3}$ are all even; or they are all odd; or $\mathbb{P}_{0}$, $\mathbb{P}_{3}$ are both even and $\mathbb{P}_{1}$, $\mathbb{P}_{2}$ are both odd; or $\mathbb{P}_{0}$, $\mathbb{P}_{3}$ are both odd and $\mathbb{P}_{1}$, $\mathbb{P}_{2}$ are both even. Each of the four cases occurs for infinitely many mixed graphs. In particular, $|\mathbb{P}_{0}|+|\mathbb{P}_{1}|+|\mathbb{P}_{2}|+|\mathbb{P}_{3}|$ is even.
We do not know any simpler proof of the Berge-Dirac Theorem than the above combination. Adding a new vertex to a mixed graph $G$ from one of the four cases and adding edges directed from the new vertex to all vertices of $G$ results in a new mixed graph from the same case. Thus it is enough to exhibit just one mixed graph for each case. An undirected edge is a graph of case 1; a directed edge is a graph of case 2; a graph on three vertices with one undirected edge, one directed edge and one missing edge is a graph of case 3; finally a graph on five vertices consisting of a directed 5-cycle and three undirected cords forming a path of length 3 and two missing cords from the same vertex is a graph of case 4.
Note that in the Berge-Dirac formula we may replace any + by a -, for example
$|\mathbb{P}_{0}|-|\mathbb{P}_{1}|-|\mathbb{P}_{2}|+|\mathbb{P}_{3}|$ is always even